情報通信システム設計特論 2 単位 (選択)

Communication Systems  高田篤 · 教授/ システム創生工学専攻電気電子創生工学コース電気電子システム講座

大家隆弘 · 教授/ システム創生工学専攻電気電子創生工学コース電気電子システム講座

【授業目的】通信方式の設計方法および多ノード間の通信制御方式の設計方法と

管理方式を理解させ，情報通信システムを設計できる能力を養う．【授業概要】通信の要素技術である線形フィルタを用いたタイミング検出及び位

相や周波数の再生，ディジタル変復調理論，およびディジタル通信方式として FDMA， TDMA， CDMA 等の無線通信および光通信ネットワーク，コンピュータネットワーク等の有線通信の多元接続方式の基礎，設計，管理技術について講述する．コンピュータネットワークでのデータ伝送プロトコル， DNS 等のネットワーク管理技術について講義を行う．またネットワークアプリケーションの実装様式について講述する． (ポートフォーリオ形式)

【授業形式】ポートフォーリオ【キーワード】phase locked loop，光伝送システム，光増幅中継伝送システム，

フォトニックネットワークシステム，コンピュータネットワーク，ネットワークアーキテクチャ，ネットワーク管理

【先行科目】『ディジタル伝送工学特論』 (1.0)，『通信工学特論』 (1.0) 【関連科目】『光半導体デバイス特論』 (0.5) 【履修要件】学部および博士前期課程における通信工学を理解していること. 【履修上の注意】 [注意] 【到達目標】

1. 無線通信 (FDMA, TDMA, CDMA など) の多ノード間通信について与えられた条件を元に通信方式を設計できる．

2. 中継システムをもつ広域データ伝送網を設計できる．

3. 有線通信 (コンピュータネットワーク) の多ノード間通信において与えられ

た要求を満たしうるアプリケーションを設計できる．【授業計画】

1. タイミング検出と位相再生

2. PLL と構成素子

3. 線形応答

4. ディジタル変復調理論

5. 光伝送システム設計

6. 光増幅中継伝送システム設計

7. フォトニックネットワークシステム

8. 移動体通信方式 (FDMA, TDMA, CDMA など)

9. コンピュータネットワークのデータ伝送プロトコル (Ethernet, ATM)

10. コンピュータネットワークのデータ伝送プロトコル (TCP, UDP)

11. ネットワーク管理技術 (DNS)

12. ネットワーク管理技術 (SNMP, MIB)

13. ネットワークアプリケーションの実装様式

14. アプリケーションプロトコルの設計

15. 分散データベースシステム

16. 分散システムの設計

【成績評価基準】レポートと質疑応答で評価．合格 60%

【教科書】 ⋄ 光増幅器とぞの応用 ⋄ 非線形光ファイバ光学

【参考書】 [参考資料]

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216701 【対象学生】開講コース学生のみ履修可能【連絡先】

⇒ 高田(E棟C3， 656-7465, takada@ee) (オフィスアワー: 高田:Tue. 13:30-1

4:30, Thu. 16:30-17:30) ⇒ 大家 (E 棟 3 階北 C-1, 088-656-7479, alex@ee.tokushima-u.ac.jp) ���� (オ

フィスアワー: 火 16:20∼ 17:20，木 16:50∼ 17:50)

Communication Systems 2 units (selection)

Atsushi Takada · Professor / Electrical and Electronic Systems, Electrical and Electronic Engineering, Systems Innovation Engineering

, Takahiro Oie · Professor / Electrical and Electronic Systems, Electrical and Electronic Engineering, Systems Innovation Engineering

Target⟩ Understanding the designing and administrating scheme of wireless communication system and multi-nodes networks. And bring up faculty of designing original telecommunication system.

Outline⟩ Timing detection and regeneration of a phase and a frequency using a linear filter, design of a digital PLL, digital modulation and demodulation, and multiple access technique, e. g. FDMA, TDMA, CDMA in wireless communications. Design of data transmission protocol over optical data transmission system and computer networks, network administration techniques, e. g. DNS, and implementation of application program for internetworking.(Portfolio style)

Style⟩ Portfolio �������⟩ phase locked loop, optical transmission system, optical transmission

system using repeater-amplifier, photonic network system, computer networks, network architecture, network management

Fundamental Lecture⟩ “Advanced Theory of Digital Transmission”(1.0), “Ad-

vanced Theory of Electrical Communication”(1.0) Relational Lecture⟩ “Photonic Semiconductor Device Physics”(0.5) Requirement⟩ Students are required to have a good understanding of communication

engineering and related subjects up to master-level. Notice⟩ [注意]

Goal⟩

1. Able to design the multiple-access wireless communication system (e.g. FDMA, TDMA, CDMA).

2. Able to design the wide-area data transmission network with repeaters.

3. Able to design the network architecture which satisfies the specified demands

on multi-nodes network (e.g. computer networks).

Schedule⟩

1. Timing detection and regeneration of a phase

2. Phase-locked loops and its components

3. Response to linear frequency variation

4. Digital modulation and demodulation

5. Designing the optical transmission system

6. Designing the optical transmission system using repeater-amplifier

7. Photonic network system 8. Mobile communication system (FDMA, TDMA, CDMA)

9. Transmission protocol of computer networks (Ethernet, ATM)

10. Transmission protocol of computer networks (TCP, UDP)

11. Network administration technique (DNS)

12. Network administration technique (SNMP, MIB)

13. Implementation of network application

14. Designing the application protocol

15. Distributed database system

16. Designing the distributed system

Evaluation Criteria⟩ Reports and presentation 100 %. More than 60 % is required

to pass the class.

Textbook⟩

⋄ Optical amplifiers and their applications ⋄ Nonlinear fiber optics

Reference⟩ [参考資料]

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216701

Student⟩ Able to be taken by only specified class(es)

Contact⟩

⇒ A. Takada (Bldg. E-C3,656-7465, takada@ee) (Office Hour: A. Takada:

Tue. 13:30-14:30, Thu. 16:30-17:30) ⇒ Oie (E-3F-C-1, +81-88-656-7479, alex@ee.tokushima-u.ac.jp) ���� (Office

Hour: Tuesday 16:20∼ 17:20, Thursday 16:50∼ 17:50)